# thanks: ORCID ID: 0009-0003-9679-035Xthanks: ORCID ID: $0$-0001-7989-9271

### v1

@@ 4 Partial bifurcation study and complex forms of drop motility
Figure 5: (a) Partial bifurcation diagram of resting and moving droplets as a function of the chemical potential $\mu_{2}$ at $D=0.45$ with the L2-norm $||h||_{2}\!=\!\left(\int h^{2}\text{d}x\right)^{1/2}$ as solution measure (period-average of $||h||_{2}$ for time-periodic states). Linearly stable states [unstable states] are represented as solid lines [dashed lines]. Only the most relevant bifurcations and branches are shown. At small $\mu_{2}$ the base state (black line) is linearly stable. It is then rendered unstable in a series of bifurcations DP1, TC1 (cf. magnification in Fig. 6). The simple traveling droplets bifurcate at a drift-pitchfork bifurcation DP3. The emerging states (red line) are unstable, but a linearly stable section is limited by two Hopf bifurcations H-, H+ (red diamonds). The insets schematically show exemplary droplet and surface tension profiles. The gray region is magnified in the bottom panel where colored regions illustrate what kind of drop motion is obtained in time simulations initialized with a flat film. (b) Analogous bifurcation diagram for $D\!=0.505$. In the magnified region only shuttling is observed in time simulations. The small inset in the bottom panel shows the temporal period $\mathcal{T}$ for both branches in the marked region. (c) Two-parameter bifurcation diagram in the $(\mu_{2},D)$-plane with $D\!=D_{1}\!=\!D_{2}$. Red lines mark the Hopf bifurcations H-, H+ that cross at C±. In the red region, the traveling states are linearly stable. The lines G and S correspond to the gluing bifurcation and the Shilnikov bifurcation. Black points mark data obtained from time simulations. The regions T and A represent possible termination points. The inset shows a magnification near the suspected T-point. Above DP1 (dash-dotted line) the base state is unstable. The two vertical dotted lines indicate the parameters of the bifurcation diagrams shown in (a) and (b). (d) Velocity of the traveling states in (a). The remaining parameters for all panels are $\mu_{1}=-1.4,r=0.3,\beta_{1}=2,\beta_{2}=0.01,\delta=1,W=10$ with a mean film thickness of $\bar{h}=7$. The computational domain is $[0,100]$ with periodic boundaries.
Figure 6: (a) Magnification of the bifurcation structure of Fig. 5(a) near the first instabilities of the base branch, i.e., near DP1, TC1 and TC2. Panel (b) shows a further magnification of the region marked in (a). The base state is first rendered unstable via the subcritical drift-pitchfork bifurcation DP1. The emerging branch of traveling states (red line) connects at a supercritical drift-pitchfork DP3 to the branch that bifurcates backwards at TC1. The number of unstable eigenvalues (EV) for each branch is given (--: linearly stable, $+$: one unstable EV, +++++ +: two unstable EVs).
@@ 4.1 Onset of self-propulsion
To explore the bifurcations that ultimately result in complex types of drop motility, we employ numerical continuation [67, 68] using the package pde2path [69] (see Appendix E). We first investigate bifurcations from the branch of resting droplets that are uniformly covered by surfactants. At $\mu_{1}\!=\!\mu_{2}$, this state corresponds to thermodynamic equilibrium. We additionally assume equal diffusion of both surfactants with $D_{1}\!=\!D_{2}\!=\!D$. A typical bifurcation diagram with control parameter $\mu_{2}$ at fixed $\mu_{1}\neq 0$ is presented in Fig. 5(a). Note that we restrict ourselves to states that are relevant to what is observed in time simulations and therefore only show a small selection of all existing branches and bifurcations. As $\mu_{2}$ is increased, the branch of uniformly covered drops is rendered unstable in a combination of a drift-pitchfork bifurcation DP1 and a transcritical bifurcation TC1 that occur in immediate succession. [^3] Note that in Ref. [55] this structure was erroneously identified as a higher-multiplicity pitchfork bifurcation. The fully resolved bifurcation structure near DP1 and TC1 is shown in Fig. 6. First, a branch of parity-broken unstable traveling states emerges subcritically at DP1. This branch then connects supercritically at a drift-pitchfork DP2 to an unstable branch of resting symmetric states that bifurcates from the base branch at TC1 and connects to it again at another transcritical bifurcation TC2 at a larger value of $\mu_{2}$. This series of bifurcations comprising DP1, TC1, DP2 and TC2 essentially produces two branches of unstable symmetric resting states that continue to exist far away from the bifurcations. These states feature either one or two surface-tension peaks near the droplet center [Fig. 5(a)]. At larger driving, a branch of traveling states emerges at a subcritical drift-pitchfork DP3 with a linearly stable section that is limited by two subcritical Hopf bifurcations H- and H+ [Figs. 5(a) and 5(d)]. This section of linearly stable states corresponds to the simple self-propelled droplets described above in Sec. 3.444In Fig. 5(a) this linearly stable section is comparatively small. However, it can be broadened significantly (up to an order of magnitude) by decreasing the diffusion constant $D$ as indicated in Fig. 5(c). The corresponding branch features several saddle-node bifurcations which are, however, only visible as “kinks” in Figs. 5(b) and 5(d). We remark that the drift-pitchfork DP3 occurs beyond a Hopf bifurcation of the base branch (not shown) that corresponds to the Hopf instability of the flat film [Fig. 2(c)]. Nevertheless, the linearly stable section of the branch of self-propelled droplets lies fully within the parameter region where the flat film is only spinodally unstable.
@@ 4.2 Crawling and shuttling droplets
We next focus on the parameter region near the primary instability of the base branch where also the simple traveling states occur [gray area in Fig. 5(a)]. In this region, we further observe strikingly complex forms of self-propulsion that we study using direct numerical simulations. First, we consider the parameter region near the destabilizing Hopf bifurcation H+. We typically find two forms of droplet motion (Fig. 7). On the one hand, droplets “crawl” across the substrate by periodically forming a single surface tension peak near the droplet center that is subsequently advected to the advancing contact line region [Figs. 7(a) and 7(b), Supplemental Video 2]. This results in phases of uniform motion interrupted by abrupt stops. On the other hand, droplets “shuttle” between two points on the homogeneous substrate by reversing their direction of travel after stopping [Figs. 7(c) and 7(d), Supplemental Video 3]. For both types of motion, we find that the fluid flow at times long after a stop (times $t_{1},t_{2},t^{\prime}_{1},t^{\prime}_{3}$ in Fig. 7) is nearly identical to the simple traveling states shown in Fig. 2. Perturbations of this flow then lead to an explosive generation of a new surface tension peak near the droplet center due to the positive feedback loop outlined in Sec. 3. Whether the droplet reverses its direction is determined by the exact position of the newly generated surface tension peak. When it is created slightly on the side of the droplet center that is opposite to the already existing peak the drop motion changes direction, otherwise it retains its direction of travel. We also find that during the advection of the new surface tension peak to the advancing contact line region, Eqs. (15) are still valid when $-\partial_{z}j_{1}$ is taken into account (Supplemental Video 4). This implies that during this phase, the local reactive dynamics is fast as compared to hydrodynamic transport.
@@
Figure 8: (a) Representation of crawling and shuttling droplets as periodic orbits in the plane spanned by the right and left contact angles $\theta_{r}$ and $\theta_{l}$. The two simple traveling states and the one-peak resting state are marked as red points and gray points, respectively. The arrows represent motion along the orbit and insets zoom into the region near the traveling states. For crawling droplets, left-crawling (strong lines) and right-crawling droplets (weak lines) are shown. They are related by reflection. With increasing $\mu_{2}$, the crawling orbits move closer to the resting state. At a critical driving $\mu_{2}^{G}$ they form a “two-winged” shuttling orbit via a gluing bifurcation. The shuttling orbits closely pass both traveling states. The driving forces are $\mu_{2}=4.140,4.1405,4.141,4.1416,4.147$ and $\mu_{2}^{G}\approx 4.1413$. The diffusion constants are $D_{1}=D_{2}=0.45$. (b) Series of shuttling orbits that successively “wind around” the traveling states. The driving forces are $\mu_{2}=4.112,4.115,4.1184$ with $D_{1}=D_{2}=0.45$. (c) Shuttling orbit at $\mu_{2}=4.14325$ and $D_{1}=D_{2}=0.503$ that closely passes by the resting state and also winds around both traveling states. (d) and (e) show height and surface tension profiles of the symmetric resting and the traveling states shown as fixed points in (a)-(c). The symmetric state is always unstable. The streamlines represent the velocity of the bulk liquid in the laboratory frame. (f), (g) Space-time representations of shuttling droplets that (f) are close to the gluing bifurcation and therefore rest longer and (g) form many loops around the traveling states and thus travel for long phases. The driving forces are (f) $\mu_{2}=4.14136$ and (g) $\mu_{2}=4.1183$ with $D_{1}=D_{2}=0.45$. The remaining parameters for all panels are as in Fig. 7. See also Supplemental Videos 5 and 6.
@@
Both crawling and shuttling motion can be understood as a periodic transition between an unstable resting and an unstable moving conformation [Figs. 8(d) and 8(e)]. The moving conformation of the drop closely resembles the (unstable) simple traveling states. [^5] Note that there exist parameter regions, notably near H+, where either crawling or shuttling and the simple traveling state are multistable [Fig. 5(a)]. In this case, it is the unstable (quasi)-time periodic state that emerges subcritically at the nearby Hopf bifurcation H+ that prevents the dynamics from converging to the simple traveling state. The resting conformation corresponds to the symmetric resting droplet with one surface tension peak. This state is always unstable [Fig. 5(a)] and any arbitrarily small breaking of the left-right symmetry induces droplet motion. Crawling and shuttling motion can both be conveniently represented as periodic orbits in a reduced phase space where the orbits closely pass the resting and the moving states. Therefore, we refer to both forms of motion as periodic (the period corresponds to the time of a single orbit roundtrip). We use a projection onto the ($\theta_{l},\theta_{r}$)-plane, where $\theta_{l}$ and $\theta_{r}$ are the left and right contact angles, respectively. They are determined from the slopes at the inflection points of $h$ (Appendix E). Examples are presented in Fig. 8(a). For crawling droplets, the orbits may be divided into three phases that represent different stages of the motion. First, the trajectory passes by either the right- ($\theta_{l}>\theta_{r}$) or the left-traveling droplet state ($\theta_{r}>\theta_{l}$). This corresponds to a right- or left-crawling droplet, respectively. Second, the trajectory is quickly expelled from the region near the traveling state and approaches the symmetric resting state. In this phase, a new surface tension peak is generated near the droplet center, the droplet abruptly stops and the dynamics slows down. During this time span the orbit also crosses the diagonal $\theta_{l}=\theta_{r}$, i.e., the advancing contact angle becomes greater than the receding one as the trajectory approaches the resting state. This does not correspond to a direction reversal but reflects a change of the droplet shape as the new surface tension peak appears. Third, as the newly generated peak is advected to the advancing contact line the trajectory departs from the unstable resting state and again approaches the moving state.
@@
Figure 10: (a) Scaling of the temporal period $\mathcal{T}$ of crawling and shuttling motion as the gluing bifurcation in is approached. The period corresponds to a single orbit roundtrip in the ($\theta_{l},\theta_{r}$)-plane. Black points denote points measured in time simulations. The solid lines show logarithmic fits with $\mathcal{T}=-A\ln|\mu_{2}-\mu_{2}^{G}|+B$ where $A,B$ and $\mu_{2}^{G}$ are fit parameters. In particular, $\mu_{2}^{G}\approx 4.14133$. (b) Periods of the first few pairs of loops that are formed in a Shilnikov(-like) mechanism for shuttling droplets. Only the stable parts of the branches are shown. From one branch to the next, $\mathcal{T}$ differs by approximately $4\pi/\omega$. The total number of loops is denoted by $l$. The other parameters are as in Fig. 7.
@@
[…]

### v3

@@ 4 Partial bifurcation study and complex forms of drop motility
Figure 5: (a) Partial bifurcation diagram of resting and moving droplets as a function of the chemical potential $\mu_{2}$ at $D=0.45$ with the L2-norm $h||_{2}\!=\!\left(\int h^{2}\text{d}x\right)^{1/2}$ as solution measure (period-average of $h||_{2}$ for time-periodic states). Linearly stable states [unstable states] are represented as solid lines [dashed lines]. Only the most relevant bifurcations and branches are shown. At small $\mu_{2}$ the resting-drop base state (black line) is linearly stable. It is then rendered unstable in a series of bifurcations DP1, TC1 (cf. magnification in Fig. 6). The simple traveling droplets bifurcate at a drift-pitchfork bifurcation DP3. The emerging states (red line) are unstable, but a linearly stable section is limited by two Hopf bifurcations H-, H+ (red diamonds). The insets schematically show exemplar droplet and surface tension profiles. The gray region is magnified in the bottom panel where colored regions illustrate what kind of drop motion is obtained in time simulations initialized with a flat film. (b) Analogous bifurcation diagram for D=0.505D\!=0.505. In the magnified region only shuttling is observed in time simulations. The small inset in the bottom panel shows the temporal period $\mathcal{T}$ for both branches in the marked region. (c) Two-parameter bifurcation diagram in the $(\mu_{2},D)$-plane with $D\!=D_{1}\!=\!D_{2}$. Red lines mark the Hopf bifurcations H-, H+ that cross at C±. In the red region, the traveling states are linearly stable. The lines G and S correspond to the gluing bifurcation and the Shilnikov bifurcation. Black points mark data obtained from time simulations. The regions T and A represent possible termination points. The inset shows a magnification near the suspected T-point. Above DP1 (dash-dotted line) the base state is unstable. The two vertical dotted lines indicate the parameters of the bifurcation diagrams shown in (a) and (b). (d) Velocity of the traveling states in (a). The remaining parameters for all panels are $\mu_{1}=-1.4,r=0.3,\beta_{1}=2,\beta_{2}=0.01,\delta=1,W=10$ with a mean film thickness of $\bar{h}=7$. The computational domain is $[0,100]$ with periodic boundaries.
Figure 6: (a) Magnification of the bifurcation structure of Fig. 5(a) near the first instabilities of the base branch, i.e., near DP1, TC1 and TC2. Panel (b) shows a further magnification of the region marked in (a). The base state is first rendered unstable via the subcritical drift-pitchfork bifurcation DP1. The emerging branch of traveling states (red line) connects at a supercritical drift-pitchfork DP2 to the branch that bifurcates backwards at the transcritical bifurcation TC1. The number of unstable eigenvalues (EV) for each branch is given (-: linearly stable, ++: one unstable EV, ++++: two unstable EVs).
@@ 4.1 Onset of self-propulsion
To explore the bifurcations that ultimately result in complex types of drop motility, we employ numerical continuation [77, 78, 79] using the package pde2path [80] (see Appendix E). We first investigate bifurcations from the base branch of resting droplets that are uniformly covered by surfactants. At $\mu_{1}\!=\!\mu_{2}$, this state corresponds to thermodynamic equilibrium. We additionally assume equal diffusion of both surfactants with $D_{1}\!=\!D_{2}\!=\!D$. A typical bifurcation diagram with control parameter $\mu_{2}$ at fixed $\mu_{1}\neq 0$ is presented in Fig. 5(a). Note that we restrict ourselves to states that are relevant to what is observed in time simulations and therefore only show a small selection of all existing branches and bifurcations. As $\mu_{2}$ is increased, the branch of uniformly covered drops is rendered unstable in a combination of a drift-pitchfork bifurcation DP1 and a transcritical bifurcation TC1 that occur in immediate succession. [^3] Note that in Ref. [55] this structure was erroneously identified as a higher-multiplicity pitchfork bifurcation. The fully resolved bifurcation structure near DP1 and TC1 is shown in Fig. 6. First, a branch of parity-broken unstable traveling states emerges subcritically at DP1. This branch then connects supercritically at a drift-pitchfork DP2 to an unstable branch of resting symmetric states that bifurcates from the base branch at TC1 and connects to it again at another transcritical bifurcation TC2 at a larger value of $\mu_{2}$. This series of bifurcations comprising DP1, TC1, DP2 and TC2 essentially produces two branches of unstable symmetric resting states that continue to exist far away from the bifurcations. These states feature either one or two surface-tension peaks near the droplet center [Fig. 5(a)]. At larger driving, a branch of traveling states emerges at a subcritical drift-pitchfork DP3 with a linearly stable section that is limited by two subcritical Hopf bifurcations H- and H+ [Figs. 5(a) and 5(d)]. Note that H- and H+ are unrelated to the Hopf bifurcations of the local reactor given by condition (13). This section of linearly stable states corresponds to the simple self-propelled droplets described above in Sec. 3.444In Fig. 5(a) this linearly stable section is comparatively small. However, it can be broadened significantly (up to an order of magnitude) by decreasing the diffusion constant $D$ as indicated in Fig. 5(c). The corresponding branch features several saddle-node bifurcations which are, however, only visible as “kinks” in Figs. 5(b) and 5(d). We remark that the drift-pitchfork DP3 occurs beyond a Hopf bifurcation of the base branch (not shown) that corresponds to the first Hopf instability of the flat film [Fig. 2(c)] (i.e., the first Hopf instability of the local reactor). Nevertheless, the linearly stable section of the branch of self-propelled droplets lies fully within the parameter region where the flat film is only spinodally unstable.
@@ 4.2 Crawling and shuttling droplets
We next focus on the parameter region near the primary instability of the base branch where also the simple traveling states occur [gray area in Fig. 5(a)]. In this region, we further observe strikingly complex forms of self-propulsion that we study using direct numerical simulations. First, we consider the parameter region near the destabilizing Hopf bifurcation H+. We typically find two forms of droplet motion (Fig. 7). On the one hand, droplets “crawl” across the substrate by periodically forming a single surface tension peak near the droplet center that is subsequently advected to the advancing contact line region [Figs. 7(a) and 7(b), Supplemental Video 2]. This results in phases of uniform motion interrupted by abrupt stops. On the other hand, droplets “shuttle” between two points on the homogeneous substrate by reversing their direction of travel after stopping [Figs. 7(c) and 7(d), Supplemental Video 3]. For both types of motion, we find that the fluid flow at times long after a stop (times $t_{0},t_{2},t^{\prime}_{1},t^{\prime}_{3}$ in Fig. 7) is nearly identical to the simple traveling states shown in Fig. 2. Perturbations of this flow then lead to an explosive generation of a new surface tension peak near the droplet center due to the positive feedback loop outlined in Sec. 3. Whether the droplet reverses its direction is determined by the exact position of the newly generated surface tension peak. When it is created slightly on the side of the droplet center that is opposite to the already existing peak the drop motion changes direction, otherwise it retains its direction of travel. We also find that during the advection of the new surface tension peak to the advancing contact line region, Eqs. (15) are still valid when $\partial_{z}j_{1}$ is taken into account (Supplemental Video 4). This implies that during this phase, the local reactive dynamics is fast as compared to hydrodynamic transport.
@@
Figure 8: (a) Representation of crawling and shuttling droplets as periodic orbits in the plane spanned by the right and left contact angles $\theta_{r}$ and $\theta_{l}$. The two simple traveling states and the one-peak resting state are marked as red points and gray points, respectively. The arrows represent motion along the orbit and insets zoom into the region near the traveling states. For crawling droplets, right-crawling (strong lines) and left-crawling droplets (weak lines) are shown. They are related by reflection. With increasing $\mu_{2}$, the crawling orbits move closer to the resting state. At a critical driving $\mu_{2}^{G}$ they form a “two-winged” shuttling orbit via a gluing bifurcation. The shuttling orbits closely pass both traveling states. The driving forces are $\mu_{2}=4.140,4.1405,4.141,4.1416,4.147$ and $\mu_{2}^{G}\approx 4.1413$. The diffusion constants are $D_{1}=D_{2}=0.45$. (b) Series of shuttling orbits that successively “wind around” the traveling states. The driving forces are $\mu_{2}=4.112,4.115,4.1181$ with $D_{1}=D_{2}=0.45$. (c) Shuttling orbit at $\mu_{2}=4.14325$ and $D_{1}=D_{2}=0.503$ that closely passes by the resting state and also winds around both traveling states. (d) and (e) show height and surface tension profiles of the symmetric resting and the traveling states shown as fixed points in (a)-(c). The symmetric state is always unstable. The streamlines represent the velocity of the bulk liquid in the laboratory frame. (f), (g) Space-time representations of shuttling droplets that (f) are close to the gluing bifurcation and therefore rest longer and (g) form many loops around the traveling states and thus travel for long phases. The driving forces are (f) $\mu_{2}=4.14136$ and (g) $\mu_{2}=4.1183$ with $D_{1}=D_{2}=0.45$. The remaining parameters for all panels are as in Fig. 7. See also Supplemental Videos 5 and 6.
@@
Both crawling and shuttling motion can be understood as a periodic transition between an unstable resting and an unstable moving conformation [Figs. 8(d) and 8(e)]. The moving conformation of the drop closely resembles the (unstable) simple traveling states. [^5] Note that there exist parameter regions, notably near H+, where either crawling or shuttling and the simple traveling state are multistable [Fig. 5(a)]. In this case, it is the unstable (quasi)-time periodic state that emerges subcritically at the nearby Hopf bifurcation H+ that prevents the dynamics from converging to the simple traveling state. The resting conformation corresponds to the symmetric resting droplet with one surface tension peak. This state is always unstable [Fig. 5(a)] and any arbitrarily small breaking of the left-right symmetry induces droplet motion. Crawling and shuttling motion can both be conveniently represented as periodic orbits in a reduced phase space where the orbits closely pass the resting and the moving states. Therefore, we refer to both forms of motion as periodic (the period corresponds to the time of a single orbit roundtrip). We use a projection onto the ($\theta_{l},\theta_{r}$)-plane, where $\theta_{l}$ and $\theta_{r}$ are the left and right contact angles, respectively. They are determined from the slopes at the inflection points of $h$ (Appendix E). Examples are presented in Fig. 8(a). For crawling droplets, the orbits may be divided into three phases that represent different stages of the motion. First, the trajectory passes by either the right- ($\theta_{l}>\theta_{r}$) or the left-traveling droplet state ($\theta_{r}>\theta_{l}$). This corresponds to a right- or left-crawling droplet, respectively. Second, the trajectory is quickly expelled from the region near the traveling state and approaches the symmetric resting state ($\theta_{l}=\theta_{r}$). In this phase, a new surface tension peak is generated near the droplet center, the droplet abruptly stops and the dynamics slows down. As the new surface tension peak appears near the droplet center, the orbit also crosses the diagonal $\theta_{l}=\theta_{r}$, i.e., the advancing contact angle becomes greater than the receding one as the trajectory approaches the resting state. This does not correspond to a direction reversal but reflects a change of the droplet shape as the new surface tension peak appears. Third, as the newly generated peak is advected to the advancing contact line the trajectory departs from the unstable resting state and again approaches the moving state.
@@
Figure 10: (a) Scaling of the temporal period $\mathcal{T}$ of crawling and shuttling motion as the gluing bifurcation is approached. The period corresponds to a single orbit roundtrip in the ($\theta_{l},\theta_{r}$)-plane. Black points denote points measured in time simulations. The solid lines show logarithmic fits with $\mathcal{T}=-A\ln|\mu_{2}-\mu_{2}^{G}|+B$ where $A,B$ and $\mu_{2}^{G}$ are fit parameters. In particular, $\mu_{2}^{G}\approx 4.14133$. (b) Periods of the first few pairs of loops that are formed in a Shilnikov(-like) mechanism for shuttling droplets. Only the stable parts of the branches are shown. From one branch to the next, $\mathcal{T}$ differs by approximately $4\pi/\omega$. The total number of loops is denoted by $l$. The remaining parameters are as in Fig. 7.
@@
Shuttling states are also found near the stabilizing Hopf bifurcation H- [Fig. 5(a)]. However, they display strikingly different behavior when the driving is increased. A typical sequence of trajectories is presented in Fig. 8(b). As $\mu_{2}$ is increased, the orbit “winds around” both traveling states, successively forming more loops (the total number of loops is denoted by $l$). Droplets then maintain their direction of travel for a longer time as the number of loops increases while both contact angles slightly oscillate. This is in stark contrast to the situation shown in Fig. 8(a), where droplets rest longer as the gluing bifurcation is approached. These two scenarios are also compared in Figs. 8(f) and 8(g) (Supplemental Videos 5 and 6). We remark that particularly in cases like in Fig. 8(g), the droplet is not restricted by the domain boundaries, i.e., it may also turn around after crossing the periodic boundary. We find that with each additional pair of loops around the traveling states, the temporal period of a single shuttle (one “back-and-forth”) increases approximately by $4\pi/\omega$ (each loop contributes $2\pi/\omega$), where $\lambda_{1,2}=\rho\pm i\omega$ with $\rho>0$ are the eigenvalues associated with the Hopf instability H- (they are the only unstable eigenvalues). The leading stable eigenvalue $\lambda_{3}<0$ is real with $|\rho/\lambda_{3}|<1$. [^6] This is a direct consequence of the nearby Hopf instability H-, where $\rho=0$. Note that $\omega$ remains approximately constant over the parameter region where we observe shuttling. We show the shuttling period over $\mu_{2}$ in Fig. 10(b) for the first few pairs of loops. We expect that the stable part of each branch is limited by two saddle-node bifurcations as the period does not diverge at the ends of the branches. They may also be limited, e.g, by period doubling bifurcations on one side as in Ref. [89].